OpenStudy (anonymous):

A block of material has a volume of 2m^3. It has a mass of 1 x 10^4 kg. Calculate the density of the material sample and express your results in g/cm^3 with appropriate scientific notation.

OpenStudy (joannablackwelder):

Density=mass/volume. You will need to convert the volume into cm^3 as well.

OpenStudy (anonymous):

So, the mass is 1 x 10^4 kg. If you don't understand scientific notation, 1 x 10^4 = 10,000, and the volume is 2m^3, or 2 cubic meters. Since density = mass / volume, and you want your answer to be in units of grams per cubic cm, you will need to divide 10,000 kg by 2 cubic meters, then convert the units from kg / m^3 to grams per cubic cm. 10,000/2 = 5,000. Your units are kg / m^3 (kilograms per cubic meter). But you want grams / cubic cm. So, \[\frac{ 10,000 kg }{ 2m^{3} } \times \frac{ 1000 g }{ 1 kg} \times \frac{ 1 m^3 }{ 1,000,000 cm^3 } = \frac{ 5 g }{ cm^3 }\]